Suppose the resistance is 1.47 Ω at a temperature of 20°C. Find the resistance at 0 degree Celsius.


1
Expert's answer
2022-03-11T09:04:02-0500

"RT=R_r\begin{bmatrix}
 1+\alpha(T-T_r)\
 
\end{bmatrix}"


Where "R_T=" Resistance of temperature T

"R_r=" Resistance of reference Temperature "T_r"

"\alpha=" Temperature coefficient of resistance at reference temperature "T_r"


"R_T=1.47[1+\alpha(0-20)]"


"=1.47(1-20\alpha)"